The market supply curves and market demand curves for books are given as follows:
Supply curve: P = 0.000002Q Demand curve: P = 11 – 0.00002Q
The short-run marginal cost curve: MC = 0.1 + 0.0009Q

At the above short-run equilibrium level, the firm is …
1. making a profit of R1 000 000
2. making a loss of R1 000 000
3. making zero economic profit
1
Expert's answer
2019-03-08T05:43:30-0500

Supply curve: P = 0.000002Q 

Demand curve: P = 11 – 0.00002Q

MC = 0.1 + 0.0009Q

Profit is maximized, when MR = MC.

MR = TR' = (P*Q)' = 11 - 0.00004Q,

11 - 0.00004Q = 0.1 + 0.0009Q,

0.00094Q = 10.9,

Q = 11596 units.

At this quantity P = 11 - 0.00002*11596 = R10.77.

If we suppose, that MC = ATC, then the total profit is:

TP = (P - ATC)*Q = (10.77 - 0.1 - 0.0009*11596)*11596 = R2708.82.

But if P = ATC, then the firm is making zero economic profit.